An athletic club has 225 feet of fencing to enclose a tennis court. What quadratic function can be used to find the maximum area of the tennis court? Find the maximum area and the lengths of the resulting fence.
Answer by lwsshak3(11628)

An athletic club has 225 feet of fencing to enclose a tennis court. What quadratic function can be used to find the maximum area of the tennis court? Find the maximum area and the lengths of the resulting fence.
**
let x=length
let y=width
2x+2y=perimeter=225
2y=225-2x
y=(225-2x)/2=112.5-x
area=x*y=x(112.5-x)=(112.5x-x^2)
area=-x^2+112.5x
complete the square
area=-(x^2+112.5x+3164.06)+3164.06
area=-(x+56.25)^2+3164.06
ans:
area≈3164 sq ft
length≈56.25 ft
width=112.5-56.25≈56.25 ft